INVESTIGATION OF THE ν_{4} and 2ν_{2} bands of ^{14}NH_{3}

C. COTTAZ, \underline{I. KLEINER}, N. LEGAY, Laboratoire de Photophysique Moléculaire, Université Paris Sud, 91405 Orsay Cédex, France; G. TARRAGO, Laboratoire de Physique Moléculaire et Applications, Université Paris Sud, 91405 Orsay Cédex, France; L. R. BROWN, J. S. MARGOLIS, H. M. PICKETT and R. L. POYNTER, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, California, 91109, U. S. A.

The line positions and intensities of ammonia are being measured and analyzed between 1200 and 2100 cm^{−1} in order to improve the molecular database in planetary applications. For this laboratory spectra are being obtained at 0.005 cm^{−1} resolution using the Fourier transform spectrometers located at Orsay and at Kitt Peak National Observatory in Arizona. The observed data contain transitions of the ν_{4} fundamental band near 1626.3 and 1627.4 cm^{−1} (for s and a inversion states respectively) and the 2ν_{2} overtone band near 1597.5 and 1882.1 cm^{−1} (for s and a inversion states respectively). These measurements are being modelled using the theoretical approach applied previsously to 3ν_{2} and ν_{2}+ν_{4} bands of 14NH_{3} in the 4 micron region^{a}. Preliminary results will be shown. Part of this research reported in this paper are perfomed at the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.

aI. Kleiner, G. Tarrago and L. R. Brown, J. Mol. Spectrosc., 173, 120 (1995)